[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO BiWeekly Status Report

    Compton Observatory Science Report #178, Friday March 6, 1995
      Chris Shrader, Compton Observatory Science Support Center
         Questions or comments can be sent to the CGRO SSC.
          Phone:    301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                          Cycle-5 Proposals

Just in case you had forgotten - Cycle-5 proposals are due one month
from tomorrow! You can still request paper copies of the NRA and
appendices by contacting the SSC (e.g. replying to this message), or
download electronic copies via ftp (on grossc.gsfc.nasa.gov, under
the nra/cycle5 directory via the Web at

Also, keep in mind that electronic proposal form submissions required
for Cycle 5. Look on the Web at the above URL, or send a blank e-mail
message to rps@cossc.gsfc.nasa.gov for instructions. 

                         Instrument Reports


EGRET operations were normal this biweekly period.  Delivery of data
to the GRO SSC remains on schedule.  Interaction with guests
investigators continues at a good level.

As noted two week ago, careful studies are in progress on the whole
gamma ray sky as observed by EGRET including all the refinements that
currently exist.  The first of these papers has been submitted to a
journal; three more should be submitted over the next ten days.  
Five papers will be presented at the International Cosmic Ray
Conference, and we are planning to submit ten or eleven to the Third
Compton Symposium, including an invited summary talk.  These latter
papers should provide a good opportunity for the community to see the
breadth  of the results that have come from EGRET.

At the moment EGRET has just begun to view a region centered at
RA=90.1 and DEC=24.7; it includes the Galactic anticenter, the Crab,
and PKS 0528+134.


OSSE detector #1 failed three times on 28 February.  Recovery from
the first  failure of the day took approximately 50 minutes and was
accomplished without  much difficulty.  The second and third failures
took considerably longer  (17 hours total) to recover from.  Commands
sent to recover from the second  failure involved moving the drive to
a calibration position.  During this  motion a third failure
occurred.  These were the fifth, sixth, and seventh such  events
since launch.  The motor drive positioning process, which in normal 
operation moves the detectors every 2 minutes, takes the detector off
line when  it detects a positioning error.  Tests for pinpointing the
cause of these drive  failures are in the planning stages.

OSSE is currently in normal operations on all four detectors.  OSSE
responded to a BATSE slew trigger during viewing period 411.5 on 23
February at 02:32 UT  and mapped the region for 12 hours.  The
slewing response to BATSE burst  triggers is currently disabled for
viewing period 412 in order to conduct  engineering tests.  We will
re-enable the slewing when the test period ends.

In viewing period 411.5 (21-28 Feb), the Z-axis target is Mrk 3
(Guest Investigator P. Nandra), and the X-axis targets are QSO
2251+158 (PI team) and NGC 7172 and PKS 2155-304 (Guest Investigator
L. Bassani).

In viewing period 412 (28 Feb - 7 Mar), the Z-axis target is the Crab
Pulsar (Engineering tests - PI team), and the X-axis targets are NGC
7172 and  PKS 2155-304 (Guest Investigator L. Bassani).

Data from viewing periods 308.0, 308.3, and 308.6 were delivered to
the Compton  GRO Science Support Center archive in the last two
weeks.  The targets during  these viewing periods were the Virgo
region sky survey, the galactic plane  region near l=337 degrees, PSR
1800-21, Mrk 421, NGC 7213,  and Sco X-1 (during reboost activities).

Recent papers which have been accepted for publication include:

"OSSE Observations of 1E 1740.7-2942 in 1992 September" by G.V. Jung
        et al. (A&A)
"OSSE Observations of 3C 273" by W.N. Johnson et al. (ApJ)
"Compton Gamma-Ray Observatory Observations of the Crab Pulsar" by
        M.P. Ulmer et al. (ApJ)
"A Search for Fast Gamma Ray Pulsars with OSSE" by P. Hertz et al.
"Highlights from OSSE on the Compton Observatory" by J.D. Kurfess
        (Proc. 17th Texas Symposium)
"OSSE Observations of Gamma Ray Emission from Centaurus A"
        by R.L. Kinzer et al. (ApJ)

A mailing including these preprints is planned in about 2 weeks.


The COMPTEL instrument is performing well and continues routine
A weak cosmic gamma-ray burst was observed within the field of view
of COMPTEL on 8 February 1995 (TJD 9756).  The BATSE/BACODINE trigger
intensity of this event was sufficiently weak that it did not meet
the threshold for an automated rapid-response.  From subsequent
analysis the COMPTEL burst position is in good agreement with the
best current BATSE/IPN position of (RA,DEC)=(338.29,56.33) degrees


BATSE continues to operate normally.

The following sources were detected by the BATSE pulsed source
monitor in the past two weeks: Her X-1, Cen X-3, 4U 1626-67, 2S
1417-624, OAO 1657-415,  GX 1+4, Vela X-1, and GX 301-2. 

As of March 1 BATSE has detected 1235 gamma-ray bursts out of a total
of 3341 on-board triggers in 1408 days of operation. There have been
749 triggers due to solar flares with emission above 60 keV.